OpenStudy (anonymous):

(3-2i)-(x+yi)=(2-3i)

myininaya (myininaya):

3-x-2i-yi=2-3i (3-x)-(2+y)i=2-3i 3-x=2 2+y=3

myininaya (myininaya):

x=1 y=1

myininaya (myininaya):

Checking: (3-2i)-(1+i)=2-2i-i=2-3i So we get an A+ on this